In how many ways can a committee of 5 people be chosen from 7 men and 3 women if it must contain at least one woman?
Answer by sudhanshu_kmr(1152)


There are 7 men and 3 women and we have to choose committee of 5

no. of ways to select 5 from total 10 = 10C5

now, we have to ignore those groups that have only men

no. of ways to select 5 person from only 7 men = 7C5


required no. = 10C5 - 7C5 = 252 - 21 = 231